# Improving DAPO from a Mixed-Policy Perspective

### v2

@@ 2 Theoretical Foundation
Suppose the objective function of the policy gradient algorithm $J\in\mathcal{J}_{n}$, where $\mathcal{J}_{n}$ is the class of finite-sum Lipschitz smooth functions, has $\sigma$-bounded gradients, and the importance weight $w=\pi_{\theta}/\pi_{\phi}$ is clipped to be bounded by $[\underline{w},\bar{w}]$. Let the learning rate $\alpha_{k}=\alpha=c/\sqrt{K}$, where
@@ 3.1 Method 1: Mixed Policy with Off-policy Guidance
where $f(x)=\frac{x}{x+\gamma}$ is a scaling function with a small constant $0<\gamma\ll 1$, and $\hat{r}_{j,t}(\theta,\phi)=\frac{\pi_{\theta}(v_{j,t}|q,v_{j,<t})}{\pi_{\phi}(v_{j,t}|q,v_{j,<t})}$.
@@
$J(\theta)=\frac{1}{2}J_{\text{on}}(\theta)+\frac{1}{2}J_{\text{off}}(\theta).$ (5)
@@ 3.1.1 Gradient Analysis and Convergence
$\nabla_{\theta}J(\theta)=0\iff\nabla_{\theta}J_{\text{off}}(\theta)=0\iff\nabla_{\theta}J_{\text{on}}(\theta)=0.$
This result guarantees that the extremum we find by optimizing $J(\theta)$ is a valid optimum in the reinforcement learning sense.
@@ 3.1.2 Benefits of the Modification
  • Variance Reduction: As training progresses and $\pi_{\theta}$ approaches $\pi_{\phi}$, the ratio $\pi_{\theta}/\pi_{\phi}$ approaches 1. In this regime, $f^{\prime}(\pi_{\theta}/\pi_{\phi})\approx\gamma$ is small. The variance of the off-policy gradient, $\operatorname*{Var}(\nabla_{\theta}J_{\text{off}})$, becomes approximately $\gamma^{2}\operatorname*{Var}(\nabla_{\theta}J^{*})$, where $\nabla_{\theta}J^{*}$ is the traditional off-policy gradient. The scaling function $f(\cdot)$ thus stabilizes training by reducing gradient variance when the policies are close.
@@ 3.2 Method 2: Reusing Zero-Reward Samples
$J_{\text{mix}}(\theta)=\underbrace{\sum_{j,t}f(\hat{r}_{j,t}(\theta,\phi))\cdot\hat{B}_{j}}_{\text{off-policy part}}+\underbrace{\sum_{k,t}\operatorname*{CLIP}(r^{(2)}_{k,t}(\theta),\hat{C}_{k},\epsilon)}_{\text{zero-reward part}}$ (8)
@@
$J(\theta)=\frac{1}{2}J_{\text{on}}(\theta)+\frac{1}{2}J_{\text{mix}}(\theta).$ (9)
@@
Similar to the first method, the analysis holds. The gradients for all three components ($J_{\text{on}},J_{1},J_{2}$) are structurally similar up to positive scaling factors. Therefore, optimizing the composite objective $J(\theta)$ will drive the gradients of all components to zero, ensuring convergence to a theoretically sound optimum.

### v3

@@ 2 Theoretical Foundation
Suppose the objective function of the policy gradient algorithm $J\in\mathcal{J}_{n}$, where $\mathcal{J}_{n}$ is the class of finite-sum Lipschitz smooth functions, has $\sigma$-bounded gradients, and the importance weight $w=\pi_{\theta}/\pi_{\phi}$ is clipped to be bounded by $\underline{w},\bar{w}]$. Let the learning rate $\alpha_{k}=\alpha=c/\sqrt{K}$, where
@@ 3.1 Method 1: Mixed Policy with Off-policy Guidance
where $f(x)=\frac{x}{x+\gamma}$ is a scaling function with a small constant $0<\gamma\ll 1$, and be careful $\hat{r}_{j,t}(\theta,\phi)=\frac{\pi_{\theta}(v_{j,t}|q,v_{j,<t})}{\pi_{\phi}(v_{j,t}|q,v_{j,<t})}$.
@@
$J_{\text{Mix\_1}}(\theta)=\frac{1}{2}J_{\text{on}}(\theta)+\frac{1}{2}J_{\text{off}}(\theta).$ (5)
@@ 3.1.1 Gradient Analysis and Convergence
$\nabla_{\theta}J_{\text{Mix\_1}}(\theta)=0\iff\nabla_{\theta}J_{\text{off}}(\theta)=0\iff\nabla_{\theta}J_{\text{on}}(\theta)=0.$
This result guarantees that the extremum we find by optimizing $J_{\text{Mix\_1}}(\theta)$ is a valid optimum in the reinforcement learning sense.
@@ 3.1.2 Benefits of the Modification
  • Variance Reduction: As training progresses and $\pi_{\theta}$ approaches $\pi_{\phi}$, the ratio $\pi_{\theta}/\pi_{\phi}$ approaches 1. In this regime, $f^{\prime}(\pi_{\theta}/\pi_{\phi})\approx\gamma$ is small. The variance of the off-policy gradient, $\operatorname*{Var}(\nabla_{\theta}J_{\text{off}})$, becomes approximately $\gamma^{2}\operatorname*{Var}(\nabla_{\theta}J_{\text{Mix\_1}}^{*})$, where $\nabla_{\theta}J_{\text{Mix\_1}}^{*}$ is the traditional off-policy gradient. The scaling function $f(\cdot)$ thus stabilizes training by reducing gradient variance when the policies are close.
@@ 3.2 Method 2: Reusing Zero-Reward Samples
$J_{\text{mix}}(\theta)=\frac{1}{\sum_{j=1}^{N_{off}}|v_{j}|+\sum_{k=1}^{N_{zero}}|w_{k}|}\bigg{(}\underbrace{\sum_{j,t}f(\hat{r}_{j,t}(\theta,\phi))\cdot\hat{B}_{j}}_{\text{off-policy part}}+\underbrace{\sum_{k,t}\operatorname*{CLIP}(r^{(2)}_{k,t}(\theta),\hat{C}_{k},\epsilon)}_{\text{zero-reward part}}\bigg{)}$ (8)
@@
$J_{\text{Mix\_2}}(\theta)=\frac{1}{2}J_{\text{on}}(\theta)+\frac{1}{2}J_{\text{mix}}(\theta).$ (9)
@@
Similar to the first method, the analysis holds. The gradients for all three components ($J_{\text{on}},J_{1},J_{2}$) are structurally similar up to positive scaling factors. Therefore, optimizing the composite objective $J_{\text{Mix\_2}}(\theta)$ will drive the gradients of all components to zero, ensuring convergence to a theoretically sound optimum.